### v3

# UseRawInputEncoding

Abstract
At it stands, the $\Lambda CDM$ model does not anticipate the early emergence of massive galaxies mcgaugh2024 . Canonical Modified Newtonian Dynamics (MOND) seems to fail at late time solar system scale tremaine -desmond and Wide-Binary scales banik . To match data, a MOND variant needs a variable MOND acceleration $a_{0}$ which is strong at high redshift galactic scale and diminishes over redshift to far below Newtonian gravity at solar system scale at late time. We found such a candidate in a relativistic frame-work. In wong , a new single-metric solution of Einstein Gravity is found for a point mass residing in an expanding universe, which apart from the Newtonian acceleration, gives rise to an additional MOND-like acceleration in which the MOND acceleration $a_{0}$ is replaced by the cosmological acceleration $\frac{1}{2}H^{2}(z)r$. This cosmological acceleration is shown to be far below Newtonian acceleration in the solar system.
[…]
pacs:
I Introduction
Mass discrepancy problems in astrophysics and cosmology are modelled extensively by working within General Relativity (GR) with an introduction of an appropriate amount of postulated cold dark matter (CDM) particles in the Stress Momentum Tensor. At very large scale, the presence of Dark matter potential after radiation-matter equality provides a good explanation for the Cosmic Microwave Background (CMB) angular power spectrum mukhanov and constrains the dark matter density parameter as well as the present time Hubble constant planck . The constraint cosmological parameters specifies the the standard $\Lambda CDM$ model. At the matter power spectrum scale, the modelled $\sigma_{8}$ at small redshift observation begins to develop tension bohringer -einasto . At late time galactic scale, there are modelling difficulties vand -mcgaugh2018 . In fact, observations support an unknown correlation between dark matter halo and its baryonic distribution salucci . At late time solar system scale, dark matter halo density is low and observation is consistent with Baryonic matter dominance, but the observed Hubble constant is in tension with the constrained value from CMB power spectrum riess -valentino . The high redshift observation of the James Webb Space Telescope (JWST) provides the biggest challenge to the $\Lambda CDM$ paradigm, which does not anticipate the early emergence of both massive galaxies mcgaugh2024 and Supermassive Black Holes (SMBH) cho . The formation of these objects seems to require a significant increase of gravity soon after recombination beyond the standard $\Lambda CDM$ model provides. As it stands, the postulated Dark matter particles remain to be found experimentally and the theoretically viable candidates are getting fewer.
An alternative approach is to search for new non-Newtonian force law (or gravitational potential) without the need of adding CDM. The simplest such paradigm is MOND proposed by Milgrom milgrom with its inclusion of a non-Newtonian radial acceleration, for a central mass $M$,
$\ddot{r}=-\sqrt{g_{N}a_{0}},\>\>\>g_{N}=\frac{GM}{r^{2}},$ (1)
which dominates the gravitational acceleration at Newtonian acceleration $g_{N}$ where $g_{N}\ll a_{0}$. In this paradigm, the value of the acceleration scale $a_{0}$ is taken to be an universal constant. Given an averaged galactic rotation speed $V$ around a central mass $M$ and the phenomenological Tully-Fisher relation $V^{4}=GMa_{0}$ (with $\ddot{r}=-V^{2}/r$), one could arrive naturally at a heuristic model with the MOND equation Eq.(1), where the acceleration scale $a_{0}$ is not a priori fixed but is given by observations of the system under consideration. It can either be an Universal constant or be dependent on the parameters of the system under consideration.
The small observed scatter of the empirical BTFR of gas rich spiral galaxies at the scale $a_{0}=1.2\times 10^{-10}ms^{-2}$, McGaugh mcgaugh2004 -Lelli lelli , makes a strong case for constant $a_{0}$ and with this choice of $a_{0}$, the MOND paradigm becomes a canonical force law. Canonical MOND can also explain the Faber-Jackson relation (FJR) at large radius in spherical galaxies, Milgrom milgrom2 . Nonetheless, in other systems such as the Giant low surface density galaxies milgrom0 , spiral galaxies with bulge angus and Early Type galaxies swaters where the observed $a_{0}$ significantly differs from the canonical value. In dwarf galaxies, the observed MOND acceleration seems to have a very larger scatter sanders0 . In galaxy clusters, MOND acceleration $\sim 4a_{0}$ is required for the central X-ray region aguirre , whilst the outer region the observed MOND acceleration is $\sim 20a_{0}$ mcgaugh2020 . In the large mass Bright cluster Galaxies (BCG), the observed MOND acceleration is $7a_{0}-20a_{0}$ tian . In large scale structure evolution simulation, the required MOND acceleration is $\sim 0.1a_{0}$ to match observation, Nusser nusser2002 . Even then, at late time, the observed Newtonian acceleration dominance at solar system scale tremaine -desmond and at Wide-Binary scale banik suggests that canonical $a_{0}$ is too large at these scales. An immediate resolution to the challenge posed by observations is to consider a MOND variant with variable $a_{0}$. Due to the resemblance of $a_{0}$ value with the cosmological acceleration $\sim cH_{0}$, Milgrom has considered a cosmologically varying $a_{0}(z)$ milgrom2015 . Other variable MOND candidate has also been considered sabine2018 .
Unlike dark particles, the Milgram MOND program alone does not provide the dark matter potential required to produce the CMB power spectrum mcgaugh2014 and the suppression of Baryonic Acoustic Oscillation (BAO) in matter power spectrum dodelson . Since the original MOND is a phenomenological theory, the hope is in finding the underlying relativistic theory from which one could obtain better understanding of the issue of $a_{0}$ and further insight in the mass deficit problem at large scales. These theories are usually achieved by modifying GR. Some MOND variant theories that survive the gravitational wave speed observations from GW170817 are MOG moffat and the Relativistic MOND by Skordis skordis . The GRB221009A observations cao also imposes a strong constraint on the Lorenz invariance of any MOND variant model.
Our starting point is a new metric solution of General Relativity for a point mass residing in an expanding universe wong in which an additional MOND-like acceleration arises, where the canonical MOND acceleration $a_{0}$ is replaced by the cosmological acceleration $\frac{1}{2}H(z)^{2}r$. For Newtonian acceleration $g_{N}$, the effective gravitational acceleration is given by $\ddot{r}=-\sqrt{g_{N}a_{0}^{VM}}$ where $\sqrt{a_{0}^{VM}}=\sqrt{g_{N}}+\sqrt{\frac{1}{2}H^{2}r}$. We call $a_{0}^{VM}$ VMOND acceleration, This metric provides the simplest realisation of the MOND idea in a relativistic framework without modifying Einstein Gravity and therefore satisfies the observed constraints set by gravitational wave speed and Lorenz invariance. At late time $H(z)=H_{0}$ and solar system scale, the non-Newtonian acceleration is negligible and Newtonian acceleration dominates wong . This avoids the problem of canonical MOND at solar system scale.
However, at early times the non-Newtonian acceleration can be dominant at some large scale, the obvious question is whether this new model can account for the early emergence of massive galaxy and produces the observed $a_{0}$ in galaxies at much later time.
In this work we examine the late time values $a_{0}^{VM}$ in non-spiral galaxies such as Ellipticals and Dwarf Spheroidal galaxies. Late time Spherical galaxy is modelled as a hydrodynamical equilibrium system in which gravitational acceleration balances the pressure within the sphere and measuring its velocity dispersions at different radius will reveal the underlying gravitational acceleration. The characteristic scale is the effective (half mass) radius $r_{e}$. At very small distances within an effective radius of Ellipticals, Bolton et. al. bolton obtains a confirmation of the fundamental plane relation (constant velocity disperson $\sigma_{r}$) which can be modelled using the ”equilibrium” Jean equation of an isothermal sphere under Newtonian acceleration where $\sigma_{r}$ is constant.
To model the observed Velocity Dispersion-Radius relation, Sanders sanders2000 introduces a small anisotropy in the form of a polytropic sphere at equilibrium, with pressure $P$, density $\rho$ and constant $K$ such that $P=K\rho^{1+\frac{1}{n}}$ with $\sigma_{r}^{2}=K\rho^{1/n}$ under the MOND acceleration. A decreasing mass density $\rho(r)$ in radial distances will then lead to a fine-tunable decrease in $\sigma_{r}$. In this polytrope model, the Faber Jackson relation
$\sigma_{r}^{4}=A(n,\beta)GM(r)a_{0}$ (2)
has a structure constant $A(n,\beta)$ which is a function of $n$ and the anisotropic parameter $\beta$ and it is found that $1613<n<$ 16 is required to fit observations. Sanders sanders2000 also highlights a degeneracy that one can also fit a data set of mass-velocity dispersion relations of spherical galaxies in Newtonian plus Dark Matter (DM) form
$\sigma_{r}^{2}=B(n,\beta)\frac{GM(r)}{r_{e}}$ (3)
where $B(n,\beta)$ includes an effective dark matter factor. This degeneracy at large radius is addressed by Sanders sanders2021 who tracks a run of velocity dispersions in the low mass density (deep MOND) region of low surface density spheroidal galaxies. The results are favourable towards MOND.
More comprehensively, in Durazo et. al. durazo , for over a large range of mass scales, the Velocity Dispersion-Radius relation show that Ellpitical galaxies typically have a Newtonian central region characterised by an effective (half-mass) radius scale $r_{e}$ above which the velocity dispersions gradually become flat and follow a non-Newtonian Faber Jackson Relation (FJR). Similar observation for a small number of elliptical galaxies is found earlier by Milgrom and Sander milgrom4 .
To model the evolution of an elliptical galaxy, Sanders sanders2007 studies a spherical overdensity evolves monolithically under MOND using simulation. It finds that small radius expanding mass shell will turnaround earlier than the large radius mass shells. This is due to an implicit comoving scale in the overdensity. Despite without a well defined MOND potential, virialisation seems plausible within a characteristic radius $r_{e}$ and the measured structure constant $A(n,z)$ in FJR seems compatible with values in the theoretical work sanders2000 .
Theoretical understanding of the virialisation process of an overdensity into an equilibrium sphere is non-trivial. While a mass shell of an overdensity cloud after turnaround in Newtonian gravity can oscillate between its largest radius and its smallest radius and follows phase mixing with neighbouring shells, the relaxation time is found to be longer than the Hubble time binney . Lydden Bell lyddenbell proposes that a time dependent potential could produce a ”violent” relaxation with relaxation time that is comparable to the crossing time. This violent relaxation in Newtonian gravity is seen in simulation by van Albada vanalbada where non uniform initial condition is required. For a spherical overdensity under MOND acceleration, Nipoti et al. nipoti show that the ”violent” relaxation can also occur with a relaxation time which is slightly longer than the Newtonian relaxation time. This same study finds that the Newtonian Fundamental plane relation requires an initial power-law matter distribution.
*In this work, we follow a spherical overdensity which monolithically evolves from recombination to virialisation in VMOND, assuming that the violent relaxation of the mass shells works in the same way as described in the simulation in MOND. In section 2, we recapitulate the essential features of our model for completeness. In section 3, we consider the turnaround and collapses of the overdensity mass shells and estimate the time for the central core to reach a time independent QSS state. In Section 4, we calculate the virialised potential and estimate the late time value of $a_{0}^{VM}$ in the outer region for a few galaxies in different environment. Section 5 is summary and discussion.
II 2: The model
For a central point mass in an expanding background with Hubble parameter $H(z)$ at redshift $z$, we find a new metric that differs from the Einstein-Straus vacuole solution, having the form wong
$ds^{2}=c^{2}d\tau^{2}-\frac{2GMa^{3}}{c^{2}r}d\varrho^{2}-r^{2}d\Omega^{2}.$ (4)
In terms of Tolman-Lemaitre formulation, the underlying free fall velocity is given by
$\dot{r}=H(z)r-\sqrt{\frac{2GM}{r}},$ (5)
where $G$, $M$ and $r$ are the Newton’s constant, the central point mass and the radial distance respectively. This equation depicts that the particle will follow the Hubble law at large distances, but at small distances it will follow a Newtonian free falling velocity. From Eq.(5), the acceleration equation takes the form
$\frac{d\dot{r}}{dt}=\dot{r}\bigg{(}\sqrt{\frac{GM}{2r^{3}}}+H\bigg{)}+r\dot{H},$ (6)
$\ddot{r}=-\frac{GM}{r^{2}}-\sqrt{\frac{H^{2}r}{2}}\sqrt{\frac{GM}{r^{2}}}+\frac{\ddot{a}}{a}r,\>\>\>\>\>\>\frac{\ddot{a}}{a}=-\frac{1}{2}H_{m}^{2}+\frac{c^{2}\Lambda}{3},$ (7)
where $H_{m}$ is the Hubble parameter in a matter only universe. Here we see that a non-Newtonian MOND-like acceleration arises due to the free fall relation Eq.(5). The effective gravitational acceleration without the cosmic background is
$\ddot{r}=-\sqrt{g_{N}a_{0}^{VM}},\>\>\>\sqrt{a_{0}^{VM}}=\sqrt{\frac{GM}{r^{2}}}+\sqrt{\frac{H^{2}r}{2}}.$ (8)
We call $a_{0}^{VM}$ the VMOND acceleration. The gravitational potential (the second term on the R.H.S we call the MOND-like potential) now takes the form
$V=\frac{GM}{r}-H\sqrt{2GMr}+\frac{1}{2}H^{2}r^{2}.$ (9)
For slow speeds, the energy equation is given by
$E=\frac{1}{2}\dot{r}^{2}+\frac{h^{2}}{2r^{2}}-V\bigg{(}H(z),r\bigg{)}.$ (10)
where $E$ is the total energy of the particle, $h$ is the angular momentum per unit mass. At the turnaround distance from central mass, we have
$r_{ta}=\bigg{(}\frac{2GM}{H^{2}}\bigg{)}^{1/3}$ (11)
Although we argue elsewhere that a different choice for $H(z)$ when there are two fluids which dominate in their own asymptotic epochs, would provide the dark matter potential required by the observed CMB angular power spectrum wong3 and matter power spectrum wong4 , as the overdensity evolution primarily occurs in the matter dominant epoch, we take the Hubble parameter at redshift $z$ to be the Friedmann-Lemaitre-Robertson-Walker (FLRW) choice
$H^{2}(z)=\frac{8\pi G}{3}\rho_{H}=\frac{8\pi G}{3}\bigg{(}\rho_{r}+\rho_{m}+\rho_{\Lambda}\bigg{)}=H_{0}^{2}\bigg{(}\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}\bigg{)},\>\>\>H^{2}_{0}=\frac{8\pi G}{3}\rho_{c},$ (12)
where for $i=r,m,\Lambda$, $\rho_{i}$ are cosmological background densities of radiation, pressureless matter which we assume is essentially baryonic matter and dark energy respectively. $H_{0}$ and $\rho_{c}$ are the Hubble parameter and the critical density of the present epoch. For $\rho_{i,0}$ denotes the density of $i$ compoent at the present epoch, we have the density parameters $\Omega_{i}=\frac{\rho_{i,0}}{\rho_{c}}$. Also the acceleration equation takes the form
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}\bigg{(}2\rho_{r}+\rho_{m}\bigg{)}+\frac{8\pi G}{3}\rho_{\Lambda}.$ (13)
For a local mass cloud in an expanding background, its central baryonic density $\rho_{b}(r)$ is given in terms of the overdensity $\delta(r)$ as
$\delta(r)\rho_{m}=\rho_{b}(r).$ (14)
In comoving coordinates, one can show that wong , the evolution equation for $\delta$ in Newtonian gravity is replaced by $\Delta(\delta)=\delta+\delta^{1/2}$ such that the overdensity evolution equation becomes
$\ddot{\Delta}+2H\dot{\Delta}=4\pi G\Delta\rho_{m},$ (15)
which has the growth solution $\Delta=\delta+\delta^{1/2}\propto a(t)$. Therefore at small $\delta\ll 1$, $\delta^{1/2}$ dominates the function $\Delta(\delta)$, so that $\delta$ grows as $\delta\propto a(t)^{2}$ which is similar to the effect of MOND. As $\delta\rightarrow 1$, the overdensity begins to grow more slowly as $\delta\propto a(t)$.
II.1 3.1: Over-density evolution from VMOND
The question therefore is whether this faster overdensity evolution in our paradigm is enough to give agreement with data. We shall suggest that it can be, but the argument is not straightforward. We have to make assumptions about galaxy formation, power spectra, simulations and energy redistribution that take us a long way beyond our original model, although they depend crucially upon its details.
From an initial overdensity $\delta_{int}$ at recombination, we can use $\delta+\delta^{1/2}\propto a(t)$ to calculate the overdensity $\delta$ at redshift $z<1080$ by
$\delta+\sqrt{\delta}=\bigg{(}\delta_{int}+\sqrt{\delta_{int}}\bigg{)}\bigg{(}\frac{1081}{1+z}\bigg{)}=\frac{A_{0}}{1+z},$ (16)
[…]
The baseline CMB average temperature variation $\frac{\delta T}{T}=1\times 10^{-5}$ corresponds to an initial baryon overdensity $\delta_{int}=3\times 10^{-5}$ and $A_{0}=5.94$. To account for high redshift large galaxies, one needs an overdensity to turnaround at sufficiently higher redshift. In sanders2007 , in a spherical galaxy formation under MOND, for galaxy mass $10^{11}M_{\odot}$ Sanders takes $\delta_{int}=1.8\times 10^{-3}$ which corresponds to $A_{0}=47.8$. [To better illustrate the effect of Eq.(16), we choose a slightly higher value $\delta_{int}=2.8\times 10^{-3}$ ($A_{0}=60$), although Sander’s choice is sufficient for most of our purposes.] We obtain the turnaround redshift $z_{ta}$ where $\delta=1$ from Eq.(16),
$291+z_{ta}=\frac{A_{0}}{2}=30,\>\>\>z_{ta}=$ 29 . (17)
[…]
When compared to the Newtonian only evolution where $\Delta=\delta$, we note that
$2.021+z_{ta}=1081\delta_{int}=3.02,\>\>\>z_{ta}=$ 2.02 . (18)
We can see the VMOND potential significantly lifts $z_{ta}$ to a much higher redshift, which could provide the model basis for the observed massive galaxies at high redshifts.
II.2 3.1: A particle’s Journey time after turnaround
From Eq.(17), we can estimate the turnaround time when the overdensity reaches $\delta=1$. In a matter only universe, we have the scale factor $a(t)\propto t^{2/3}$, so that
$H(z)=\frac{\dot{a}}{a}=\frac{2}{3t}.$ (19)
At redshift $z_{ta}$ the turnaound time is
$t_{ta}=\frac{2}{3H(z_{ta})}.$ (20)
After $\delta\geq 1$, we can calculate the particle free fall time as follows. The $E=h=0$ energy equation of a point around a central mass $M$ is Eq.(5)
$\dot{r}=Hr-\sqrt{\frac{2GM}{r}}=\sqrt{\frac{2GM}{r}}\bigg{(}\sqrt{\frac{H^{2}r^{3}}{2GM}}-1\bigg{)}.$ (21)
In the matter dominant epoch, $H^{2}=\frac{4}{9t^{2}}$ (c.f. Eq.(19) ). Setting
$y=\frac{2}{3}\frac{r^{3/2}}{\sqrt{2GM}},\>\>\>y=xt,$ (22)
Eq.(21) simplifies to
$\frac{dy}{dt}=\frac{y}{t}-1,\>\;\>\bigg{(}\frac{dx}{dt}=-\frac{1}{t}\bigg{)}.$ (23)
Since $x$ is a dimensionless variable, one obtains a solution with a constant $t_{0}$ to be fixed.
$\frac{y}{t}=x=\ln\bigg{(}\frac{t_{0}}{t}\bigg{)}.$ (24)
To fix $t_{0}$, we use the turnaround time $t=t_{ta}$ where $\dot{r}=0\>(\dot{y}=0)$. Then, from Eq.(23)
$\frac{dy}{dt}=0,\>\>\>\frac{y(r_{ta})}{t_{ta}}=1=\ln\bigg{(}\frac{t_{0}}{t_{ta}}\bigg{)},$ (25)
so that $t_{0}=et_{ta}$. From Eq.(24) as $r\rightarrow 0$ (we have $y\rightarrow 0$) where $t$ goes from $t_{ta}$ to reach its closest approach at $t_{1}=et_{ta}$. Here the free fall time for a turnaround particle is $t_{ff}=(e-1)t_{ta}=1.72t_{ta}$.
In Binney-Tremaine (BT) binney (pp-379-382), after turnaround, the evolution of particles on the mass shells is described by the Poisson equation and the collisionless Boltzmann equation, in which the density of phase points in a small volume around the phase point of any oscillating particle is constant. Different initial energies of a particle will stretch the original phase space patch into a filament. When the filaments are sufficiently (phase-)mixed the resulting macroscopic distribution function will tend to its equilibrium state.
Most importantly, Lynden-Bell (LB) lyddenbell argues that a star’s potential should vary (violently) in space and time. Its varying total energy $E$ then leads to a redistribution of energies between stars and provides a much faster relaxation, called ”Violent relaxation”. More specifically, LB argues that in this case, the solution of the system consisting of collisionless-Boltzmann equations involves intermingled filaments on smaller and smaller scales. For macroscopic observables, LB introducees a coarse-grained distribution function which is a local average of distribution function over the filaments. This coarse-grained distribution is expected to reach maximum entropy on a short dynamical time $t_{D}$. In the case of stellar systems, this coarse-grained distribution will tend to the Maxwell-Boltzmann distribution. Works on the violent relaxation mechanism are still ongoing. Recently, the violent relaxation is observed in the laboratory lovi . The collapse via phase mixing and violent relaxation, is called virialisation (binney pp734). After the process is complete, in a spherical collapse, the final sphere should satisfy the Virial Theorem, which we will use in the next subsection.
We note that different authors use different definitions of dynamical time unit to do simulations or for analytic estimates. For example, in BT binney pp.268, the (Newtonian) free fall time
$t_{ff}=\frac{1}{4}\sqrt{\frac{3\pi}{2G\rho_{m}}}$ (26)
is used. Lynden-Bell lyddenbell uses the dynamical time
$t_{D}=\frac{3}{4}\sqrt{\frac{1}{2\pi G\rho_{m}}}.$ (27)
We follow the definition in Nipoti et al. nipoti , ciotti ,
$t_{dyn}=2t_{N},\>\>\>t_{N}=\sqrt{\frac{r^{3}}{GM}},$ (28)
mainly because they simulate the spherical gravitational collapse of both Newtonian and MOND system, so that it provides a simpler reference time-scale for collapse in VMOND, which should be at the same order.
In lyddenbell , the estimated violent relaxation time is $t_{D}\sim 0.91t_{N}$, therefore, a coarse-grain equilibrium central sphere can appear at $t\leq t_{ta}+t_{N}$. In a Newtonian simulation in binney (pp 383), where a different free fall time $t_{ff}\sim 1.11t_{N}$ is used, after $t=t_{ta}+1.61t_{N}$, a tight minimum configuration has developed. By $t=t_{ta}+4.8t_{N}$, the core is settled to a Quasi-Stationary State (QSS) (a not fully equilibrium but time independent state). This is attributed to the stretching and folding of filaments debuyl . At $t=t_{ta}+9t_{N}$, an equilibrium sphere is developed. However, it does not preclude the possibility that some outer large mass shells remain outsdie of the virialised sphere.
This is a common feature seen in gravitational collapse simulation works, vanalbada , nipoti , sanders2007 , ciotti . In general, a tight central region can form after $t=t_{ta}+t_{N}$. A QSS can be reached at $t=t_{ta}+4t_{N}$, ciotti , and a complete virialisation is reached at $t=t_{ta}+8t_{N}$, nipoti . The gravitational collapse simulations in nipoti and ciotti are done for both Newtonian and MOND acceleration.
II.3 3:3: Can a Stable Central Region form before $z>6.5$
We note that JWST currently only provides the image of a large and bright central region for the very high redshift $z>11$ galaxies and only the galaxy morphologies for the $6.51.5<z<$ 6.5. It does not provide enough information on how far the galaxy virialisation has progressed. The tight central configuration obtained after $t_{N}$ could arguably provide a reasonable description of the observed bright central region in high redshift galaxies.
At $\delta=1$, the overdensity mass inside $r_{ta}$ is given by $M=\frac{4\pi}{3}\rho(z_{ta})r_{ta}^{3}$, so that from Eq.(28) $t_{dyn}$ takes the value
$t_{dyn}=2\sqrt{\frac{r_{ta}^{3}}{GM}}=\frac{2\sqrt{2}}{H(z_{ta})}.$ (29)
The time to virialisation $t_{vr}$ is given by
$t_{vr}=t_{ta}+t_{ff}+Nt_{dyn}=\frac{2}{3H(z_{ta})}\bigg{(}e+3\sqrt{2}N\bigg{)}=\frac{2}{3H(z_{vr})},$ (30)
where $z_{vr}$ is the virialisation redshift and $N$ is the number of $t_{dyn}$ after $t_{ff}$ required to reach an effective virialisation. From Eq.(30), we obtain
$1+z_{vr}=\frac{1+z_{ta}}{(e+3\sqrt{2}N)^{2/3}}$ (31)
Since in VMOND, the $t_{ff}=1.72t_{ta}=0.814t_{N}$ is shorter than $t_{D}\sim 0.91t_{N}$ in the violent relaxationlyddenbell . The faster free fall time implies that the mixing would be more violent and we assume that after $t_{ff}$ ($N=0$), a tight minimum configuration could occur. Eq.(31) shows that this occurs at $z_{vr}=14.46$.
For a Quasi-Stationary-State (QSS) obtained in simulation binney , ciotti , on taking $N=1$, QSS occurs at
$z_{vr}=\frac{1+z_{ta}}{3.64}-1=7.24.$ (32)
An effectively complete virialisation occurs at $N=3$ nipoti , where
$z_{vr}=\frac{1+z_{ta}}{6.2}-1=3.81.$ (33)
Our initial overdensity $\delta_{int}=2.8\times 10^{-3}$ is a moderate choice within the Barkana-Loeb constraint. The analysis above suggests that with this $\delta_{int}$, the VMOND potential is sufficient to provide for a tight minimum galactic configuration at redshifts higher than $z_{vr}\sim 14.46$, and a stable (Quasi-Stationary State) galactic core by $z\geq 7$ before the central core evolves through the range $6.51.5<z<$ 6.5, as observed in ferreira2 .
III 4: The velocity disperson in a virialised VMOND potential
When a mass shell turnarounds and decouples from its background, the energy of the mass shell is in principle conserved during its subsequent motions. However, during violent relaxation and virialisation, some of the total energy of the large number of mass shells could be lost due to physical reasons, such as particle ejections, conduction and radiation, before the system comes close to equilibrium. In an already virialised system, the change of a particle’s potential in time is given by the change of energy (see Lynden-Bell lyddenbell )
$\frac{dE}{dt}=\frac{\partial\Phi}{\partial t},$ (34)
which, after equilibrium, would mainly be due to radiation. This progresses slowly and when it does lose energy, the velocity dispersions could remain stable.
Consider the central region of a protogalactic cloud that relaxes through a violent relaxation under the VMOND acceleration and reaches a meta-equilibrium at the order of a few dynamical times. Given the time average kinetic energy $K=\frac{1}{2}Mv^{2}$, where $v$ is the average speed, and the average potential V, the virialised relation is $02K+V=$ 0. At this point, (the average velocity $v$ no longer resembles the initial averaged velocity) but is constrained by the average potential $V$.
We assume the mass density of the cloud at turnaround $\delta=1$ (distance $r$) is given by
$\rho_{m}(r)=\rho_{m}\bigg{(}\frac{r_{i}}{r}\bigg{)}^{S}=\bar{\rho}r^{-S},\>\>\>\bar{\rho}=\rho_{m}r_{i}^{S}.$ (35)
where $S>0$ is some constant. The central mass upto a shell at $r$ is
[…]
The virialised potential binney is given by
[…]
For the Newtonian acceleration, using $M(r)$ in Eq.(36)
[…]
$=$ $-G\frac{(4\pi\bar{\rho})^{2}}{(3-S)}\frac{R^{5-2S}}{(5-2S)}=-c_{1}(S)\frac{GM^{2}}{R};\>\>\>c_{1}(S)=\bigg{(}\frac{3-S}{5-2S}\bigg{)}.$
where $M=M(R)=4\pi\bar{\rho}R^{3-S}/(3-S)$. For an uniform density, $S=0$, one recovers the virial potential $V=-\frac{3}{5}\frac{GM^{2}}{R}$ (Binney binney , pp 64). For $S=2$, one obtains the Newtonian potential energy $V=-\frac{GM^{2}}{R}$. The calculation above suggests that the virialised potential Eq.(37) provides a reasonable description of the effective potential of the virialised object.
For the VMOND acceleration, assuming for the time being that there is ”no potential energy loss” during virialisation, the virialised potential is given by
[…]
$\Delta V(R)=-\frac{H(z_{ta})}{\sqrt{2}}4\pi\bar{\rho}\sqrt{\frac{4\pi G\bar{\rho}}{(3-S)}}\frac{R^{5-\frac{3}{2}S}}{5-\frac{3}{2}S}$ (40)
That is,
$\Delta V(R)=-MH(z_{ta})\sqrt{\frac{GMR}{2}}c_{2}(S),\>\>\>\>c_{2}(S)=\bigg{(}\frac{3-S}{5-\frac{3}{2}S}\bigg{)}.$ (41)
For a stationary orbit where the kinetic energy is dominated by an averaged $3-D$ velocity dispersion $v^{2}=\sigma^{2}=3\sigma_{r}^{2}$, ($\sigma_{r}$ is the velocity dispersion in the radial direction,) taking also a particle’s turnaround radius $r_{ta}^{3}=\frac{2GM}{H^{2}(z)}$, we have
$\sigma^{2}=c_{1}(S)\frac{GM}{R}\bigg{(}1+\frac{c_{2}(S)}{c_{1}(S)}\sqrt{\frac{H^{2}(z_{ta})R^{3}}{2GM}}\bigg{)}=c_{1}(S)\frac{GM}{R}\bigg{(}1+\frac{c_{2}(S)}{c_{1}(S)}\sqrt{\frac{\rho(z_{ta})}{\rho_{b}(R)}}\bigg{)}.$ (42)
where $M=\frac{4\pi}{3}\rho(R)R^{3}$. With an emphasis on the long distance $R$ factor, we have
$\sigma^{4}=c_{2}^{2}(S)\bigg{(}\frac{1}{2}H^{2}(z_{ta})R\bigg{)}\bigg{(}1+\frac{c_{1}(S)}{c_{2}(S)}\sqrt{\frac{2GM}{H^{2}(z_{ta})R^{3}}}\bigg{)}^{2}.$ (43)
At turnaround where $\delta=1$, $r=r_{ta}$ at $z=z_{ta}$, the overdensity mass
$M=\frac{4\pi}{3}r_{ta}^{3}\rho(z_{ta}).$ (44)
After the overdensity decouples from the background contracts to radius $r<R$. To evaluate the virial potential at $R$, it is convenient to define $n=\frac{r_{ta}}{R}$ and obtain
$\sigma^{4}=c_{2}(S)^{2}GM\bigg{(}\frac{1}{2}H(z_{ta})^{2}R\bigg{)}\bigg{(}1+\frac{c_{1}(S)}{c_{2}(S)}n^{3/2}\bigg{)}^{2}.$ (45)
From Eq.(42) and Eq.(43), one notices that there is a central small $R$ region where Newtonian acceleration dominates. There is also an outer large $R$ region where a MOND-like acceleration dominates. Between these two regions there is a transition from Newtonian dominance to MOND dominance. Qualitatively, this is the result of Durazo et al. durazo .
If there is large potential energy loss during virialisation and we have to evaluate the non-Newtonian virial potential at some other redshift $z=z_{vr}$. Eq.(41) becomes
$\Delta V(R)=-MH(z_{vr})\sqrt{\frac{GMR}{2}}c_{2}(S).$ (46)
The central mass $M$ can take a different form
$M=\frac{4\pi}{3}r_{ta}^{3}\rho(z_{ta})=\frac{4\pi}{3}r_{ta}^{3}\rho(z_{vr})\bigg{(}e+3\sqrt{2}N\bigg{)}^{2}.$ (47)
From Eq.(42), we obtain a modified Newtonian velocity dispersion
$\sigma^{2}=c_{1}(S)\frac{GM}{R}\bigg{(}1+\frac{c_{2}(S)}{c_{1}(S)}\sqrt{\frac{H^{2}(z_{vr})R^{3}}{2GM}}\bigg{)},$ (48)
or in Faber-Jackson form
$\sigma^{4}=c_{2}(S)^{2}GM\bigg{(}\frac{1}{2}H(z_{vr})^{2}R\bigg{)}\bigg{(}1+\frac{c_{1}(S)}{c_{2}(S)}n^{3/2}\sqrt{\frac{\rho(z_{ta})}{\rho(z_{vr})}}\bigg{)}^{2},\>\>\>\frac{\rho(z_{ta})}{\rho(z_{vr})}=\bigg{(}e+3\sqrt{2}N\bigg{)}^{2}.$ (49)
The last equality on the R.H.S. comes from Eq.(31).
III.1 4.1: The velocity dispersion - $a_{0}$ relation
In elliptical galaxy, the observables are the velocity dispersions $\sigma$ and central mass $M$ (through a best estimated mass to light ratio). The observed mass deficit in the $\sigma$-$M$ relation in Newtonian gravity at large distances is attributed to a dark particle halo or a new gravitational acceleration scale $a_{0}$. In practice, from the observed $\sigma$ and $M$, the value of $a_{0}$ is obtained with an assumed density profile and MOND interpolation function. For example, Milgrom milgrom uses an isothermal sphere to evaluate the large radius limit of 3-D isotropic velocity $\sigma$. The MOND acceleration $a_{0}$ is given by
$\sigma^{4}=\frac{4}{9}GMa_{0},$ (50)
However, in a study involving spherical galaxies over 7 order of magnitude in mass durazo , where more general density profile is appropriate. The following $\sigma-M$ relation is used.
$\sigma^{4}=\frac{1}{9}GMa_{0}.$ (51)
III.2 4.2 The $\sigma(R)$ profile of spherical galaxies
For elliptical galaxies, the characteristic scale is the half mass radius $R_{e}$ (the radius from within which half of the galaxy mass is contained). Durazo et al. durazo , considers spherical galaxies over 7 orders of magnitude in mass, it is found that the velocity dispersion is nearly constant within $R_{e}$, called the Fundamental plane. This constant velocity dispersions profile below $R_{e}$ can be modelled by Newtonian gravity with $\rho(r)\propto r^{-2}$ which leads to $M(R)\propto R$ and a constant $\sigma$. The velocity dispersion Newtonian behaviour is well described by the small $R$ region in Eq.(48). Beyond $R_{e}$, durazo presents the $\sigma(R)$ profile for different elliptical galaxies without providing actual masses. The $\sigma(R)$ initially decreases non-linearly and transitions to a nearly constant profile at large $R$.
To see whether Eq.(45) can lead to the observed behaviour in durazo without assuming galactic mass. We choose $S=0$ (so that $c_{1}(S)=c_{2}(S)=3/5$), whence Eq.(45) becomes, for a fixed $M$,
$\frac{\sigma^{4}}{r_{ta}}=GM\bigg{(}\frac{3}{5}\bigg{)}^{2}\bigg{(}\frac{1}{2}H_{0}^{2}\bigg{)}\Omega_{b}(1+z_{ta})^{3}\frac{1}{n}\bigg{(}1+n^{3/2}\bigg{)}^{2}.$ (52)
From Eq.(44) $r_{ta}$ is fixed given a fixed $M$ and $z_{ta}$. Eq.(52) can be simplified to
$\sigma^{4}=Constant\times\frac{1}{n}\bigg{(}1+n^{3/2}\bigg{)}^{2}.$ (53)
For small $R_{e}<R<r_{ta}$, where $n=r_{ta}/R\gg 1$, we have
$\sigma^{2}\propto n\propto\frac{1}{R}.$ (54)
which is a Newtonian behaviour. For $R$ closer to $r_{ta}$ such that $n-1=\frac{xR}{R}=\epsilon$, where $\epsilon=\frac{r_{ta}-R}{R}\ll 1$
$\sigma^{4}\propto 4\bigg{(}1+\frac{1}{2}\epsilon\bigg{)},$ (55)
which approaches to a constant when $R\rightarrow r_{ta}$. Between the Newtonian behaviour region and the nearly constant value limit, there clearly exists a transition. To see the usefulness of the above analysis, we make the crude approximation that $\sigma_{\infty}$ is evaluated at $\epsilon=0$, and obtain
$\frac{\sigma(R)}{\sigma_{\infty}}=\bigg{[}\frac{1}{4n}\bigg{(}1+n^{3/2}\bigg{)}^{2}\bigg{]}^{1/4}.$ (56)
From Fig.3 and Fig.4 in Durazo et al. durazo , we note that $\sigma(R)/\sigma_{\infty}$ descends from $2.5$ at $r\sim 0.1R_{e}$ to $1$ at $R\geq 5R_{e}$. For the system we consider below at $M_{*}=10^{10.5}M_{\odot}$ and $r_{ta}=13R_{e}$. At small $R=0.1R_{e}$ ( $n=130$), we obtain $\sigma(R)/\sigma_{\infty}\sim 2.38$. At $r=6R_{e}$, $\sigma(R)/\sigma_{\infty}=1.19$, which is in qualitative agreement with Fig.3 and Fig.4 in Durazo et al. durazo . We wish to point out that $R_{e}$ for a central mass $M$ has significant uncertainty, so that Eq.(56) can match the plots in Fig.3 and Fig.4 closely at small and large $R$ and qualitatively in the intermediate region is remarkable. To further test the validity of Eq.(45), we will calculate the $a_{0}^{VM}$ value at large distances where the non-Newtonian acceleration dominates and see whether it matches the canonical value $a_{0}$.
III.3 4.3 Value of $a_{0}^{VM}$ for large mass galaxy
Our attention so far is on the massive galaxies evolved monolithically from an appropriate overdensity at last scattering, we continue our consideration to these virialised large galaxies and see whether the $a_{0}^{VM}$ in Eq.(51) with $\sigma$ from Eq.(45), which is not a priori fixed to a canonical value, can lead to a MOND-like acceleration which can fit observations. For large distances behaviour, we choose, from suess , a large star forming elliptical galaxy with mass $M_{*}=10^{10.5}M_{\odot}$ with $R_{e}\sim 3kpc$ and data is given out to $r_{18}=6R_{e}$. We use the Milky way ”gas to star” ratio $M_{g}=(3.43/6.07)M_{*}$ wong2 . From $M=M_{*}+M_{g}$ and $\rho(M)=\rho(z_{ta})$ at $\delta=1$, However, for the present time critical density $\rho_{c}$, the spherical mass $M$ is fixed at $z_{ta}$ by
$M=\frac{4\pi}{3}\Omega_{b}\rho_{c}(1+z_{ta})^{3}r_{ta}^{3},$ (57)
where $\Omega_{b}$ is the baryon density parameter. We find that $r_{ta}=39.2kpc$. To evaluate $a_{0}^{VM}$, we make a simplification by writng phenomenological MOND acceleration $a_{0}$ in multiples of $\frac{1}{2}H_{0}^{2}r_{18}$, where $r_{18}=6R_{e}$, the contraction factor $n=\frac{r_{ta}}{6R_{e}}=2.18$.
$a_{0}=\frac{1}{6}cH_{0}=\bigg{(}\frac{1}{2}H_{0}^{2}r_{18}\bigg{)}\gamma;\>\>\gamma=\frac{c/H_{0}}{3\times 18kpc}=7.41\times 10^{4},$ (58)
From
$a_{0}^{VM}=\frac{9\sigma^{4}}{GM}=\bigg{(}\frac{1}{2}H_{0}^{2}r_{18}\bigg{)}\gamma,$ (59)
choosing $z_{ta}=29$ and $\Omega_{b}=0.05$,
$\gamma=9\bigg{(}\frac{3}{5}\bigg{)}^{2}\Omega_{b}(1+z_{ta})^{3}\bigg{(}1+n^{3/2}\bigg{)}^{2}.$ (60)
$\gamma=7.78\times 10^{4},\>\>\>a_{0}^{VM}(6R_{e})=1.05a_{0},$ (61)
in remarkable qualitative agreement with the $a_{0}$ of MOND asnalysis. In some other cases, durazo and milsanders , the velocity dispersions can be given upto to $R=7R_{e}$. At this distances, the same calculation yields $a_{0}^{VM}(7R_{e})=0.86a_{0}$. We can extrapolate the calculation to $R=r_{ta}=13R_{e}$ and obtain $a_{0}^{VM}(13R_{e})=0.51a_{0}$. At very large distances, Eq.(51) persists with a weakly decreasing $a_{0}$. In practice, like in the Radial Acceleration Relation in lelli , the observed gravitation acceleration at different mass scale presented is an averaged acceleration obtained from velocity dispersion observations over different distances from mass centre.
If we assume very significant potential energy loss during the virialisation process and the potential virialises at some $z_{vr}$, Eq.(49) leads to
$\gamma=9\bigg{(}\frac{3}{5}\bigg{)}^{2}\Omega_{b}(1+z_{vr})^{3}\bigg{(}1+\sqrt{(e+3\sqrt{2}N)^{2}n^{3}}\bigg{)}^{2}.$ (62)
If we take $N=0$, $z_{vr}=14.46$,
$\gamma=5.68\times 10^{4},\>\>\>a_{0}^{VM}(6R_{e})=0.77a_{0}.$ (63)
Choosing the QSS state, $N=1$, gives
$8.241+z_{vr}=\frac{1+z_{ta}}{3.64}=$ 8.24 . (64)
We obtain
$\gamma=4.96\times 10^{4},\>\;\>a_{0}^{VM}(6R_{e})=0.67a_{0},$ (65)
That $a_{0}^{VM}\sim a_{0}$ suggests that potential energy loss (for example ejection of particles) is small during the virialisation process, the major energy loss is from the kinetic energy (for example, particle slowing down).
Next, we consider a larger galaxy at the scale of BCG with mass $10^{11.5}M_{\odot}$, which has small $R_{e}=1-4kpc$. If we assume $z_{ta}=29$ as before, the turnaround radius is larger at $r_{ta}=2.15\times 39kpc$, we can evaluate $a_{0}^{VM}$ around $18kpc$ as before and obtain
$a_{0}^{VM}(18kpc)=7.37a_{0}.$ (66)
We also note that at $a_{0}^{VM}(15kpc)=9.94a_{0}$. In tian , for 54 MaNGA BCGs the observed MOND acceleration is $a_{0}^{VM}=7.5a_{0}-20a_{0}$.
From the observed galactic mass and a simple overdensity assumption at recombination that we can obtain in the VMOND model $a_{0}^{VM}\sim a_{0}$ (the canonical MOND acceleration), at $10^{10.5}M_{\odot}$ and $a_{0}^{VM}\sim 10a_{0}$ (the observed non-canonical MOND value) at $10^{11.5}M_{\odot}$ is remarkable.
III.4 4.4: A Brief Note on Dwarf Spherical Galaxy
Kinematics of Tidal Dwarf Galaxies (TDG) which are formed by debris at late time $z\simeq 0$ is found to be compatible with Newtonian gravity with no Dark matter lelli3 . It also posts a challenge to the canonical MOND paradigm. In the VMOND paradigm, the turnaround redshift $z_{ta}=0$. the non-Newtonain potential in Eq.(52) is at $O(10^{-3})$ times the Newtonian potential and the velocity dispersion therefore should be dominated by Newtonian acceleration dynamics.
On the other hand, low surface brightness spheroidal galaxies around the Milky Way galaxy are found to have large velocity dispersions and is compatible with MOND for some galaxies sanders2021 . Here we consider the galaxy Carina in sanders2021 where $M=2.4\times 10^{6}M_{\odot}$, the effective radius $r_{eff}=0.369\>kpc$, the line of sight velocity dispersions is averaged around $6.6\>km/s$. We assume that Carina is a small part of the large overdensity which formed the Milky Way. Instead of gravitationally collapsing to the Milky Way mass centre, this small cloud also decouples from the cosmic background at $z_{ta}=29$ at sufficiently high systematic angular momentum which keeps the gas cloud to stay in a large stable orbit around the Milky Way centre. This gas subsequently gravitationally collapses onto its own mass centre lying in the large orbit. We assume that this small cloud gravitational collapse process behaves similarly to the collapse process as the large central galaxy discussed above. For the mass of Carina, the turnaround radiius is $2.67kpc$. Consider $R=0.5kpc$, $n=5.34$
$\gamma=9\bigg{(}\frac{3}{5}\bigg{)}^{2}\bigg{(}\Omega_{b}(1+z_{ta})^{3}\bigg{(}\frac{0.5kpc}{18kpc}\bigg{)}\bigg{)}(1+n^{3/2})^{2}=2.16\times 10^{4}.$ (67)
The velocity dispersion
$\sigma(0.5kpc)=\bigg{(}\frac{1}{9}GMa_{0}^{VM}\bigg{)}^{1/4}=5.9km/s.$ (68)
This is within the line of sight observation scatter $\sigma=6.6\pm 2.6\>km/s$ in sanders2021 .
IV 5: Summary and Discussion
Any Solution to the Missing gravity problem in Astrophysics and Cosmology, either by postulating new particles or New force law (or gravitational potential), needs to be a relativistic theory and will necessarily impact all physical systems involving gravity. Both $\Lambda CDM$ and canonical MOND models are currently facing difficulties in their respective systems. For a central mass in an expanding background, our proposed solution is a new metric to General Relativity in which a variable MOND acceleration arises in its geodesic equation. This MOND variant solution comes within a relativistic framework and therefore avoids the constraints set by Gravitational wave speed and Gamma Ray observations on Lorenz invariance. The non-Newtonian MOND-like acceleration, which depends on $H(z)$ and radius, is found to fall below Newtonian gravity in our solar system and thus match observations that are problematic for canonical MOND wong . In this work we use the new model VMOND to consider elliptical galaxy formed monolithically from an overdensity at recombination. After the overdensity mass shells turnaround and free fall to the mass centre, to estimate the time it takes for the overdensity mass shells to reach its virialised state, we have to assume the gravitational collapse simulation results in the literature under Newtonian and MOND acceleration which show that a central core can form and reach a time independent Quasi Stationary State after another $2t_{N}$. For VMOND with a realistic overdensity at recombination, we find that the time to form a QSS state is $z\geq 7$, which could explain the JWST result that a large number of galaxies can evolve through $6.51.5>z>$ 6.5 without morphology change. In the virialised potential, the averaged Velocity dispersons-Radius relation matches observations in Durazo et al. durazo . Given a galaxy mass, we can calculate $a_{0}^{VM}$ using the equilibrium velocity dispersion values obtained from the galaxy virialised potential. We find that for massive galactic scale $M=10^{10.5}M_{\odot}$, $a_{0}^{VM}\sim a_{0}$ at $6R_{e}$ which is the typical observation limit. At Bright Cluster Galaxy scale $M>10^{11.5}M_{\odot}$, we find that $a_{0}^{VM}>7.3a_{0}$ which is consistent with the observed non-canonical MOND value $7.2a_{0}-20a_{0}$ for a group of BCGs. Tidal Dwarf Galaxies are formed by debris at late time. The non-Newtonian potential is evaluated at $z_{ta}=0$ which has a much smaller value than the Newtonian potential and the galaxy kinematics therefore should follow the dominant Newtonian acceleration dynamics lelli3 . From our model, we show also that it is possible to obtain observed high velocity dispersions in Dwarf Spheroidal galaxy orbiting at large radius around the Milky Way.
In Summary: Application of the VMOND model to spherical galaxy systems results in late time $a_{0}^{VM}$ value which is both non-canonical and matches observations. This remarkable result further supports the idea that VMOND can address missing mass problem upto the galactic scale. The case for spiral galaxy is reported in wong2 . The missing mass problem in CMB power spectrum and Matter power spectrum observations are addressed elsewhere wong3 -wong4 .
Acknowledgements
We thank Prof. Salucci for helpful comments on an earlier version of the text.
Data availability statement
Data Sharing not applicable to this article as no datasets were generated or analysed during the current study.
Competing Interests
The authors have no conflicts of interest to declare that are relevant to the content of this article.